Calculate the total interest earned for the investment below.
Principal = $8200
Rate of Compound Interest per Year = 2.8%
Compounding Period = quarterly
Time = 15 years

Answer by math_tutor2020(3817)


Compound interest formula
A = P*(1+r/n)^(n*t)

where,
A = final amount
P = deposit
r = annual interest rate in decimal form
n = compounding frequency
t = number of years

In this case
P = 8200
r = 0.028
n = 4
t = 15

A = P*(1+r/n)^(n*t)
A = 8200*(1+0.028/4)^(4*15)
A = 8200*(1.007)^(60)
A = 12461.8375505852
A = 12461.84
This is the final balance.

Subtract off the deposit to determine the interest earned.
i = A - P
i = 12461.84 - 8200
i = 4261.84


Answer: $4261.84